# Reducing Stochastic Games to Semidefinite Programming

### v1

@@ 3 From stochastic games to max-average constraints
In this section, we focus on the decision problem stated in Theorem 2.3 above: given a stopping simple stochastic game we want to decide whether the value of a given state is at least $1/2$. This problem can be reduced to max-average constraint satisfaction problems. The ideas for this reduction are not new, but we are not aware of an explicit reference for this fact, and we therefore present a short proof in this section. In the next section, we will show that max-average CSPs then further reduce to the feasibility problem of non-Archimedean semidefinite programs. The fact that stochastic games can be encoded as non-Archimedean SDPs was proven in [AGS18] and our reduction is very similar. While their approach relies on reducing stochastic games to ergodic games, we omit this step and obtain a slightly more straightforward reduction.
@@
An instance of a max-average constraint satisfaction problem is a conjunction of constraints, each of which has one of the following forms:
@@
The computational task is to determine whether such a conjunction has a solution over ${}\mathbb{Q}\cup\{-\infty\}$ (equivalently, over ${}\mathbb{R}\cup\{-\infty\}$).
Note that max-average CSP instances may also use constraints of the form $x_{0}\leq\min(x_{1},\dots,x_{k})$ where $k\in\mathbb{N}$, $x_{0}\leq c$, or $x_{0}\geq c$ for some rational constant $c$, because the respective relations are primitively positively definable over the other relations.
We now show that the problem of solving stochastic mean payoff games can be reduced to max-average CSPs. A reduction to slightly more general CSPs, including constraints of the form $x_{0}\leq c+x_{1}$, follows from [AGS18, Theorem 18] and is based on a Collatz–Wielandt property of order-preserving and additively homogeneous maps. However, for our purposes, it is useful that the reduction does not use these additional constraints. Instead, we give a proposition that characterizes the value of a stopping simple stochastic game. It was already proven in [Con92] that this value is a unique fixed point of an operator involving min, max, and averaging operations. By replacing the fixed point equations with inequalities, we obtain a set of certificates which prove that the value is at least $1/2$. This observation is proven, e.g., in [BBDFS23] and gives inequalities very similar to the inequalities on Shapley operators used in [AGS18].
@@
$x_{i}$ $\leq\max_{(i,j)\in E}\{x_{j}\}$ for every $\text{for every }\in V_{\mathrm{Max}},$ (1)
$x_{i}$ $\leq\min_{(i,j)\in E}\{x_{j}\}$ for every $\text{for every }i\in V_{\mathrm{Min}},$
$x_{i}$ $\leq\frac{1}{2}\sum_{(i,j)\in E}x_{j}$ for every $\text{for every }i\in V_{\mathrm{Rand}},$
@@
Proof.
Let $(Q)$ denote the system of inequalities obtained from Eq. 1 by removing the inequality $x_{k}\geq 1/2$. By [Con92] or [BBDFS23, Theorem 77], the value of the game satisfies $(Q)$ with inequalities replaced by equalities. Hence, if the value at state $k$ is at least $1/2$, then the value of the game satisfies Eq. 1. Moreover, the value of the game is a rational vector by [Con92, Lemma 2] or [BBDFS23, Corollary 18]. Conversely, if a point $x\in\mathbb{Q}_{\geq 0}^{n}$ satisfies $(Q)$, then [BBDFS23, Lemma 83] implies that for all $i\in[n]$, the entry $x_{i}$ is not greater than the value of the game at state $i$. Hence, if we further suppose that $x_{k}\geq 1/2$, then the value of the game at state $k$ is at least $1/2$. ∎
4 From max-average constraints to non-Archimedean SDPs
By “lifting” the inequalities Eq. 1 to Puiseux series, we get the reduction from max-average constraint satisfaction problems to non-Archimedean SDPs. A (formal real) Puiseux series is a series of the form
@@ 4 From max-average constraints to non-Archimedean SDPs
where $t$ is a formal parameter, the number $n$ is a positive integer, $a$ is an integer, and the coefficients $c_{i}$ are real. We further assume that $c_{0}\neq 0$ for every Puiseux series except the zero series. The number $a/n$ is called the valuation of the series $\bm{x}$ and is denoted by $\operatorname*{\mathsf{val}}(\bm{x})$, with the convention that $\operatorname*{\mathsf{val}}(0)=-\infty$. Puiseux series can be added and multiplied in the natural way. Moreover, they can be ordered by defining $\bm{x}\geq 0$ if $c_{0}\geq 0$ and $\bm{x}\geq\bm{y}$ if $\bm{x}-\bm{y}\geq 0$. It is known that Puiseux series form a real closed field, see, e.g., [BPR06, Section 2.6] for a detailed discussion. [^1] In this work, we use a convention in which one interprets $t$ as a “very large” parameter. Many authors use the opposite but equivalent convention in which $t$ is a very small parameter. This explains the differences in notation between our work and [BPR06]. We denote the field of Puiseux series by $\mathbb{K}$. For any $\bm{x},\bm{y}\in\mathbb{K}$, the following properties of the valuation function are immediate from the definitions:
@@
An instance $\phi$ with variables $x_{1},\dots,x_{n}$ of a max-average constraint satisfaction problem has a solution in ({}$(\mathbb{Q}\cup\{-\infty\})^{n}$ if and only if the following system of inequalities $S_{\phi}(t)$, which is a (quadratic) semidefinite program over Puiseux series, has a solution in $\mathbb{K}^{n}$: for every variable $x_{i}$ of $\phi$ we have a variable $\bm{x}_{i}$, and the inequalities are
@@
Proof.
Suppose that $\bm{x}=(\bm{x}_{1},\dots,\bm{x}_{n})\in\mathbb{K}^{n}$ satisfies Eq. 3. Then, $\operatorname*{\mathsf{val}}(\bm{x})$ belongs to ({}$(\mathbb{Q}\cup\{-\infty\})^{n}$ and satisfies $\phi$ by the properties of the valuation function Eq. 2. Conversely, if $(x_{1},\dots,x_{n})\in(\mathbb{Q}\cup\{-\infty\})^{n}$ satisfies $\phi$, then $(t^{x_{1}},\dots,t^{x_{n}})$ (with the convention that $t^{-\infty}=0$) satisfies Eq. 3.
This amounts to the feasibility problem for a semidefinite program: the linear (in)equalities can be expressed by diagonal matrices and each quadratic condition in combination with the non-negativity of all variables is just the positive semidefiniteness of a symmetric 2-by-2 matrix. Conjunctions of such constraints can be expressed by LMIs for block diagonal matrices. ∎
@@ 5 From non-Archimedean to real SDPs
To conclude, we aim to reduce the non-Archimedean semidefinite program from Eq. 3 to a real semidefinite program. We will do this in two steps. First, we replace the formal parameter $t$ by a real number that is sufficiently large (doubly exponential in the size of the CSP). In fact, any non-Archimedean semialgebraic feasiblity problem can be reduced to a semialgebraic feasibility problem over the reals by replacing $t$ with a sufficiently large real number. This follows from the quantifier elimination in real closed fields coupled with the definition of the order in Puiseux series. By using the effective bounds on quantifier elimination discussed, e.g., in [BPR06], we can give a bound on this “sufficiently large” number. Second, we will show that this SDP with very large coefficients can be reduced to another SDP that has coefficients of polynomial bit-size.
Before presenting the proofs, we note that a reduction from non-Archimedean SDPs coming from games to real SDPs with large coefficients was already studied in [AGS18, Theorem 36]. However, the bounds given in [AGS18] rely on an additional assumption on the underlying game. In particular, they cannot be applied directly to the SDP given in Eq. 3. For max-average systems coming from games, one could overcome this restriction by perturbing the underlying game slightly. Nevertheless, the bounds coming from quantifier elimination which we present below do not require any additional assumptions and can be of independent interest.
@@
Let $P_{1},\dots,P_{m}\in\mathbb{Z}[x_{0},x_{1},\dots,x_{n}]$ be a collection of polynomials with integer coefficients and let {,,<,>,=,}$\diamond\in\{\leq,\geq,<,>,=,\neq\}^{m}$ be a sign pattern. Consider the semialgebraic set over Puiseux series defined by
@@
Then, $\bm{S}$ is nonempty if and only if $S(L)$ is nonempty for all sufficiently large $L$. More precisely, there exists an absolute constant $C>0$ with the following property. If the maximal degree of any polynomial $P_{i}$ is bounded by $d\geq 2$ and $s$ denotes the maximal bit-size of any coefficient of $P_{i}$, then for $L>2^{sd^{Cn}}$ either all $S(L)$ are empty or all $S(L)$ are nonempty. In the former case $\bm{S}$ is empty and in the latter case $\bm{S}$ is nonempty.
Proof.
Consider the first-order formula $\phi(x_{0})$ in the language of ordered rings given by
@@
By Tarski’s quantifier elimination theorem [BPR06, Theorem 2.77], there exists a quantifier-free formula $\psi(x_{0})$ that is equivalent to $\phi(x_{0})$ in the theory of real closed fields. The bounds on quantifier elimination given in [BPR06, Theorem 14.6] show that $\psi(x_{0})$ can be taken to be of the form
@@
where $\delta_{ijk}\in\{-1,0,1\}$ and $Q_{ijk}\in\mathbb{Z}[x_{0}]$ are univariate polynomials with integer coefficients. Furthermore, every $Q_{ijk}$ has degree bounded by $d^{O(n)}$ and every coefficient of every $Q_{ijk}$ has bit-size bounded by $sd^{O(n)}$. Hence, the Cauchy bound [BPR06, Corollary 10.4] implies that every real root of $Q_{ijk}$ has absolute value bounded by $2^{sd^{O(n)}}$ (provided that $Q_{ijk}$ is a non-zero polynomial). In particular, the sign of $Q_{ijk}(L)$ is the same for all $L>2^{sd^{O(n)}}$. Therefore, either $S(L)$ are empty for all such $L$ or they are nonempty for all such $L$. Moreover, $Q_{ijk}(t)\in\mathbb{K}$ is a Puiseux series with finitely many terms. By definition, the sign of this series is the same as the sign of $Q_{ijk}(L)\in\mathbb{R}$. Hence, $\bm{S}$ is nonempty if and only if $S(L)$ is nonempty. ∎
@@
There exists an absolute integer constant $M>0$ such that an instance $\phi$ of a max-average constraint satisfaction problem has a solution if and only if the real SDP $S_{\phi}(L)$ obtained from $S_{\phi}(t)$ (see Theorem 4.1) by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$, is feasible.
Proof.
Consider the system of inequalities given in Eq. 3. Since $\bm{x}_{i}\geq 0$ for all $i$, any equality of the form $\bm{x}_{i}=t^{c_{j}}=t^{p_{j}/q_{j}}$ can be replaced with an equivalent equality $\bm{x}_{i}^{q_{j}}=t^{p_{j}}$ if $p_{j}\geq 0$ and $t^{-p_{j}}\bm{x}_{i}^{q_{j}}=1$ if $p_{j}<0$. If we further replace the parameter $t$ by a variable symbol $x_{0}$, Eq. 3 becomes a system of polynomial inequalities of degree at most $W$ and such that the coefficients of the polynomials appearing in these inequalities are integers in $\{-1,0,1\}$. Hence, by Theorem 5.1, there exists $M>0$ such that the system Eq. 3 has a solution over Puiseux series if and only if the system obtained by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$ has a solution over the real numbers. Therefore, the claim follows from Theorem 4.1. ∎
In Proposition 5.2, we obtain a (real) semidefinite program, whose size, however, is exponential in the size of $\phi$ and therefore does not provide a polynomial-time reduction to the semidefinite program feasiblity problem. However, by the following result we can also reduce to a semidefinite program of polynomial representation size.
@@
Proof.
To prove the claim, we extend the constructions of SDPs with solutions of doubly-exponential size given in [PK97, Ram97]. For $n=0$, the set $\{x\in\mathbb{R}\mid x=1\}$ is the desired spectrahedron. Suppose that $n\neq 0$ and let $|n|=\sum_{i=0}^{k}\bar{b}_{k-i}2^{i}$ be the binary representation of $|n|$ for $\bar{b}_{i}\in\{0,1\}$. Let $b_{i}=\operatorname*{\mathrm{sign}}(n)\bar{b}_{i}\in\{-1,0,1\}$ for all $i$, so that $n=\sum_{i=0}^{k}b_{k-i}2^{i}$. We first show that the set $\{x\in{\mathbb{R}}\mid x\geq 2^{n}\}$ is the projection of a spectrahedron $S$ of representation size that is polynomial in the bit-size of $n$. Let $S$ be defined by the linear matrix inequality
@@
over the spectrahedron $S$ specified above. Then, the dual is of the form $\sup\,\langle B,Y\rangle$ such that $\langle A_{i},Y\rangle=c_{i}$ for every $i\in\{0,\dots,k\}$ and $Y\succeq 0$, where $(c_{0},\dots,c_{k})=(0,\dots,0,1)$. Clearly, the primal is strictly feasible and bounded by $2^{n}$ from below, so we may use strong duality (Theorem 2.2) and obtain that $\sup\,\langle B,Y\rangle=2^{n}$ and that the supremum is attained for some $Y$. Then the following primitive positive formula over spectrahedral sets with one free variable $x_{k}$ defines the singleton relation $\{2^{n}\}$:
$\exists x_{0},\dots,x_{k-1},y_{1,1},\dots,y_{m,m}:A_{0}x_{0}+\cdots+A_{k}-B\succeq 0$
@@
where $Y$ denotes $(y_{i,j})_{i,j\in\{1,\dots,m\}}$ for $m=2(k+1)$. As we have mentioned before, spectrahedral shadows are closed under primitive positive definability, and the representation size of the resulting LMI is bounded by the representation sizes for the conjuncts. We thus obtain an LMI of representation size that is polynomial in $k$. ∎
@@
Proof.
We have already specified a first reduction to a (real) SDP in Proposition 5.2. To obtain a proper polynomial-time reduction, we modify this SDP as follows. Fix an instance $\phi$; let $c_{i}$, $p_{i}$, $q_{i}$, and $W$ be as introduced before Proposition 5.2. Define $D\coloneqq\prod_{i}q_{i}$ and $K\coloneqq 2DW^{Mn}=D2^{1+Mn\log_{2}W}$, where $M$ is the integer constant from Proposition 5.2. Furthermore, let $L\coloneqq 2^{K}$. By Proposition 5.2, $\phi$ has a solution if and only if $S_{\phi}(L)$ has a solution over the real numbers. This SDP has constraints of the form $x_{i}=L^{c_{j}}=2^{c_{j}K}$. Since $K$ is an integer divisible by $D$, the number $c_{j}K$ is an integer for all $j$. Hence, by introducing new variables, we may replace the constraints $x_{j}=2^{c_{j}K}$ by the corresponding LMIs of small representation size from Theorem 5.3. The resulting SDP can be computed in polynomial time. Furthermore, it is feasible if and only if the original SDP is feasible, which proves the statement. ∎

### v2

@@ 3 From stochastic games to max-average constraints
In this section, we focus on the decision problem stated in Theorem˜2.3 above: given a stopping simple stochastic game we want to decide whether the value of a given state is at least $1/2$. This problem can be reduced to max-average constraint satisfaction problems (max-average CSPs). The ideas for this reduction are not new, but we are not aware of an explicit reference for this fact, and we therefore present a short proof in this section. In the next section, we will show that max-average CSPs, and even the larger class of max-plus-average CSPs, then further reduce to the feasibility problem of non-Archimedean semidefinite programs. The fact that stochastic games can be encoded as non-Archimedean SDPs was proven in [6] and our reduction is very similar. While their approach relies on reducing stochastic games to ergodic games, we omit this step and obtain a slightly more straightforward reduction.
@@
An instance of a max-plus-average CSP is a conjunction of constraints, each of which has one of the following forms:
@@
The computational task is to determine whether such a conjunction has a solution over $\mathbb{Q}\cup\{-\infty\}$ (equivalently, over $\mathbb{R}\cup\{-\infty\}$). If the instance does not contain constraints of the form $2.$, we call it a max-average instance.
Note that max-average CSP instances may also use constraints of the form $x_{0}\leq\min(x_{1},\dots,x_{k})$ where $k\in\mathbb{N}$, $x_{0}\leq c$, or $x_{0}\geq c$ for some rational constant $c$, because the respective relations are primitively positively definable using max-average constraints.
We now show that the problem of solving stochastic mean payoff games can be reduced to max-average CSPs.
A reduction to max-plus-average CSPs follows from [6, Theorem 18] and is based on a Collatz–Wielandt property of order-preserving and additively homogeneous maps. We provide a more direct reduction to the smaller class of CSPs without constraints of the form $x_{0}\leq c+x_{1}$ as an intermediate step for our reduction. To do so, we give a proposition that characterizes the value of a stopping simple stochastic game.
It was already proven in [23] that this value is a unique fixed point of an operator involving min, max, and averaging operations. By replacing the fixed point equations with inequalities, we obtain a set of certificates which prove that the value is at least $1/2$. This observation is proven, e.g., in [13] and gives inequalities very similar to the inequalities on Shapley operators used in [6].
@@
$x_{i}$ $\leq\max_{(i,j)\in E}\{x_{j}\}$ $\text{for every }i\in V_{\mathrm{Max}},$ (1)
$x_{i}$ $\leq\min_{(i,j)\in E}\{x_{j}\}$ $\text{for every }i\in V_{\mathrm{Min}},$
$x_{i}$ $\leq\frac{1}{2}\sum_{(i,j)\in E}x_{j}$ $\text{for every }i\in V_{\mathrm{Rand}},$
@@
Proof 3.3.
Let $(Q)$ denote the system of inequalities obtained from Equation˜1 by removing the inequality $x_{k}\geq 1/2$. By [23] or [13, Theorem 77], the value of the game satisfies $(Q)$ with inequalities replaced by equalities. Hence, if the value at state $k$ is at least $1/2$, then the value of the game satisfies Equation˜1. Moreover, the value of the game is a rational vector by [23, Lemma 2] or [13, Corollary 18]. Conversely, if a point $x\in\mathbb{Q}_{\geq 0}^{n}$ satisfies $(Q)$, then [13, Lemma 83] implies that for all $i\in[n]$, the entry $x_{i}$ is not greater than the value of the game at state $i$. Hence, if we further suppose that $x_{k}\geq 1/2$, then the value of the game at state $k$ is at least $1/2$.
4 From max-plus-average constraints to non-Archimedean SDPs
By “lifting” the inequalities in Definition 3.1 to Puiseux series, we get the reduction from max-plus-average constraint satisfaction problems to non-Archimedean SDPs. A (formal real) Puiseux series is a series of the form
@@ 4 From max-plus-average constraints to non-Archimedean SDPs
where $t$ is a formal parameter, the number $n$ is a positive integer, $a$ is an integer, and the coefficients $c_{i}$ are real. We further assume that $c_{0}\neq 0$ for every Puiseux series except the zero series. The number $a/n$ is called the valuation of the series $\bm{x}$ and is denoted by $\operatorname*{\mathsf{val}}(\bm{x})$, with the convention that $\operatorname*{\mathsf{val}}(0)=$-$\infty$. Puiseux series can be added and multiplied in the natural way. Moreover, they can be ordered by defining $\bm{x}\geq 0$ if $c_{0}\geq 0$ and $\bm{x}\geq\bm{y}$ if $\bm{x}-\bm{y}\geq 0$. It is known that Puiseux series form a real closed field, see, e.g., [11, Section 2.6] for a detailed discussion. [^1] In this work, we use a convention in which one interprets $t$ as a “very large” parameter. Many authors use the opposite but equivalent convention in which $t$ is a very small parameter. This explains the differences in notation between our work and [11]. This field is non-Archimedean, meaning that the order does not satisfy the Archimedean property: the series $\bm{x}(t)=t$ is larger than any natural number, i.e., $t\geq m$ for all $m\in\mathbb{N}$. We denote the field of Puiseux series by $\mathbb{K}$. For any $\bm{x},\bm{y}\in\mathbb{K}$, the following properties of the valuation function are immediate from the definitions:
@@
An instance $\phi$ with variables $x_{1},\dots,x_{n}$ of a max-plus-average constraint satisfaction problem has a solution in $(\mathbb{Q}\cup\{-\infty\})^{n}$ if and only if the following system of inequalities $S_{\phi}(t)$, which is a (quadratic) semidefinite program over Puiseux series, has a solution in $\mathbb{K}^{n}$: for every variable $x_{i}$ of $\phi$ we have a variable $\bm{x}_{i}$, and the inequalities are
@@
Proof 4.2.
Suppose that $\bm{x}=(\bm{x}_{1},\dots,\bm{x}_{n})\in\mathbb{K}^{n}$ satisfies Equation˜3. Then, $\operatorname*{\mathsf{val}}(\bm{x})$ belongs to $(\mathbb{Q}\cup\{-\infty\})^{n}$ and satisfies $\phi$ by the properties of the valuation function Equation˜2. Conversely, if $(x_{1},\dots,x_{n})\in(\mathbb{Q}\cup\{-\infty\})^{n}$ satisfies $\phi$, then $(t^{x_{1}},\dots,t^{x_{n}})$ (with the convention that $t^{-\infty}=0$) satisfies Equation˜3.
This amounts to the feasibility problem for a semidefinite program: the linear (in)equalities can be expressed by diagonal matrices and each quadratic condition in combination with the non-negativity of all variables is just the positive semidefiniteness of a symmetric 2-by-2 matrix. Indeed, the symmetric matrix $\begin{pmatrix}x_{1}&x_{3}\\ x_{3}&x_{2}\end{pmatrix}$ is positive semidefinite if and only if $x_{1}\geq 0$, $x_{2}\geq 0$, and $x_{1}x_{2}\geq x_{3}^{2}$. (More generally, a symmetric matrix is positive semidefinite if and only if all of its principal minors are nonnegative.) These are precisely the inequalities that appear in Equation˜3, together with some linear inequalities. A conjunction of such inequalities can then be expressed by a single LMI given by block-diagonal matrices (a symmetric block-diagonal matrix is positive semidefinite if and only if every block is positive semidefinite).
@@ 5 From non-Archimedean to real SDPs
To conclude, we aim to reduce the non-Archimedean semidefinite program from Equation˜3 to a real semidefinite program. We will do this in two steps. First, we replace the formal parameter $t$ by a real number that is sufficiently large (doubly exponential in the size of the instance). In fact, any non-Archimedean semialgebraic feasiblity problem can be reduced to a semialgebraic feasibility problem over the reals by replacing $t$ with a sufficiently large real number. This follows from the quantifier elimination in real closed fields coupled with the definition of the order in Puiseux series. By using the effective bounds on quantifier elimination discussed, e.g., in [11], we can give a bound on this “sufficiently large” number. Second, we will show that this SDP with very large coefficients can be reduced to another SDP that has coefficients of polynomial bit-size.
Before presenting the proofs, we note that a reduction from non-Archimedean SDPs coming from games to real SDPs with large coefficients was already studied in [6, Theorem 36]. However, the bounds given in [6] rely on an additional assumption on the underlying game. In particular, they cannot be applied directly to the SDP given in Equation˜3. For max-average systems coming from games, one could overcome this restriction by perturbing the underlying game slightly. Nevertheless, the bounds coming from quantifier elimination which we present below do not require any additional assumptions and can be of independent interest. We recall that Tarski’s quantifier elimination allows to eliminate quantifiers of first-order $\tau$-formulas, with signature $\tau$ given by $\tau=\{+,\cdot,0,1,\leq\}$.
@@
Let $P_{1},\dots,P_{m}\in\mathbb{Z}[x_{0},x_{1},\dots,x_{n}]$ be a collection of polynomials with integer coefficients and let $\diamond\in\{\leq,\geq,<,>,=,\neq\}^{m}$ be a sign pattern. Consider the semialgebraic set over Puiseux series defined by
@@
Then, $\bm{S}$ is nonempty if and only if $S(L)$ is nonempty for all sufficiently large $L$. More precisely, there exists an absolute constant $C>0$ with the following property. If the maximal degree of any polynomial $P_{i}$ is bounded by $d\geq 2$ and $s$ denotes the maximal bit-size of any coefficient of $P_{i}$, then either for all $L>2^{sd^{Cn}}$ we have $S(L)=\emptyset$, or for all $L>2^{sd^{Cn}}$ we have $S(L)\neq\emptyset$. In the former case $\bm{S}$ is empty and in the latter case $\bm{S}$ is nonempty.
Proof 5.2.
Both sets $S(L)$ and $\bm{S}$ can be defined by the same first-order formula with a free variable $x_{0}$ that can be substituted by a real number $L$ (to obtain $S(L)$) or a Puiseux series $t$ (to obtain $\bm{S}$). Let $\tau=\{+,\cdot,0,1,\leq\}$ and consider the first-order $\tau$-formula $\phi(x_{0})$ given by
@@
By Tarski’s quantifier elimination theorem [11, Theorem 2.77], there exists a quantifier-free formula $\psi(x_{0})$ that is equivalent to $\phi(x_{0})$ in the theory of real closed fields. The bounds on quantifier elimination given in [11, Theorem 14.6] show that $\psi(x_{0})$ can be taken to be of the form
@@
where $\delta_{ijk}\in\{-1,0,1\}$ and $Q_{ijk}\in\mathbb{Z}[x_{0}]$ are univariate polynomials with integer coefficients. Furthermore, every $Q_{ijk}$ has degree bounded by $d^{O(n)}$ and every coefficient of every $Q_{ijk}$ has bit-size bounded by $sd^{O(n)}$. Hence, the Cauchy bound [11, Corollary 10.4] implies that every real root of $Q_{ijk}$ has absolute value bounded by $2^{sd^{O(n)}}$ (provided that $Q_{ijk}$ is a non-zero polynomial). In particular, the sign of $Q_{ijk}(L)$ is the same for all $L>2^{sd^{O(n)}}$. Therefore, either $S(L)$ is empty for all such $L$ or $S(L)$ is nonempty for all such $L$. Moreover, $Q_{ijk}(t)\in\mathbb{K}$ is a Puiseux series with finitely many terms. By definition, the sign of this series is the same as the sign of $Q_{ijk}(L)\in\mathbb{R}$. Hence, $\bm{S}$ is nonempty if and only if $S(L)$ is nonempty.
@@
There exists an absolute integer constant $M>0$ such that an instance $\phi$ of a max-plus-average constraint satisfaction problem has a solution if and only if the real SDP $S_{\phi}(L)$ obtained from $S_{\phi}(t)$ (see Theorem˜4.1) by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$, is feasible.
Proof 5.4.
Consider the system of inequalities given in Equation˜3. Since $\bm{x}_{i}\geq 0$ for all $i\in\{1,\dots,n\}$, any equality of the form $\bm{x}_{i}=t^{c_{j}}=t^{p_{j}/q_{j}}$, for an integer $q_{j}\geq 1$, can be replaced with an equivalent equality $\bm{x}_{i}^{q_{j}}=t^{p_{j}}$ if $p_{j}\geq 0$ and $t^{-p_{j}}\bm{x}_{i}^{q_{j}}=1$ if $p_{j}<0$. Likewise, any inequality of the form $\bm{x}_{i}\leq t^{p_{j}/q_{j}}\bm{x}_{l}$ can be replaced by the equivalent inequality $\bm{x}_{i}^{q_{j}}\leq t^{p_{j}}\bm{x}_{l}^{q_{j}}$ if $p_{j}\geq 0$ and $t^{-p_{j}}\bm{x}_{i}^{q_{j}}\leq\bm{x}_{l}^{q_{j}}$ if $p_{j}<0$. If we further replace the parameter $t$ by a variable symbol $x_{0}$, Equation˜3 becomes a system of polynomial inequalities of degree at most $W$ and such that the coefficients of the polynomials appearing in these inequalities are integers in $\{-1,0,1\}$. Hence, by Theorem˜5.1, there exists $M>0$ such that the system Equation˜3 has a solution over Puiseux series if and only if the system obtained by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$ has a solution over the real numbers. Therefore, the claim follows from Theorem˜4.1.
In Proposition˜5.3, we obtain a (real) semidefinite program, whose size, however, is exponential in the size of $\phi$ and therefore does not provide a polynomial-time reduction to the semidefinite program feasiblity problem. We now show how to reduce to a semidefinite program of polynomial representation size. This is based on the following result.
@@
Proof 5.6.
To prove the claim, we extend the constructions of SDPs with solutions of doubly-exponential size given in [52, 55]. For $n=0$, the set $\{x\in\mathbb{R}\mid x=1\}$ is the desired spectrahedron. Suppose that $n\neq 0$ and let $|n|=\sum_{i=0}^{k}\bar{b}_{k-i}2^{i}$ be the binary representation of $|n|$ for $\bar{b}_{i}\in\{0,1\}$. Let $b_{i}=\operatorname*{\mathrm{sign}}(n)\bar{b}_{i}\in\{-1,0,1\}$ for all $i$, so that $n=\sum_{i=0}^{k}b_{k-i}2^{i}$. We first show that the set $\{x\in{\mathbb{R}}\mid x\geq 2^{n}\}$ is the projection of a spectrahedron $S$ of representation size that is polynomial in the bit-size of $n$. Let $S$ be defined by the linear matrix inequality
@@
over the spectrahedron $S$ specified above. Then, the dual is of the form $\sup\,\langle B,Y\rangle$ such that $\langle A_{i},Y\rangle=c_{i}$ for every $i\in\{0,\dots,k\}$ and $Y\succeq 0$, where $(c_{0},\dots,c_{k})=(0,\dots,0,1)$. Clearly, the primal is strictly feasible and bounded by $2^{n}$ from below, so we may use strong duality (Theorem˜2.2) and obtain that $\sup\,\langle B,Y\rangle=2^{n}$ and that the supremum is attained for some $Y$. Then the following primitive positive formula over spectrahedral sets with one free variable $x_{k}$ defines the singleton relation $\{2^{n}\}$:
$\exists x_{0},\dots,x_{k-1},y_{1,1},\dots,y_{m,m}:A_{0}x_{0}+\cdots+A_{k}x_{k}-B\succeq 0$ (4)
@@
where $Y$ denotes $(y_{i,j})_{i,j\in\{1,\dots,m\}}$ for $m=2(k+1)$. As we have mentioned before, spectrahedral shadows are closed under primitive positive definability, and the representation size of the resulting LMI is polynomially bounded by the representation sizes for the conjuncts. Indeed, the matrices $A_{i}$ and $B$ have coefficients in $0,\pm 1/2,\pm 1,\pm 2$} and are of size $k\times k$ which is polynomial in the bit-size of $n$. Hence, the system of matrix inequalities Equation˜4 can be encoded using a number of bits that is polynomial in the bit-size of $n$. This system of inequalities can be then encoded by a single block-diagonal LMI as in the proof of Theorem˜4.1. This increases the number of necessary bits by a polynomial factor. We thus obtain an LMI of representation size that is polynomial in $k$.
@@
Proof 5.8.
We have already specified a first reduction to a (real) SDP in Proposition˜5.3. To obtain a proper polynomial-time reduction, we modify this SDP as follows. Fix a max-average instance $\phi$; let $c_{i}$, $p_{i}$, $q_{i}$, and $W$ be as introduced before Proposition˜5.3. Define $D\coloneqq\prod_{i}q_{i}$ and $K\coloneqq 2DW^{Mn}=D2^{1+Mn\log_{2}W}$, where $M$ is the integer constant from Proposition˜5.3. Furthermore, let $L\coloneqq 2^{K}$. By Proposition˜5.3, the instance $\phi$ has a solution if and only if $S_{\phi}(L)$ has a solution over the real numbers. This SDP has constraints of the form $x_{i}=L^{c_{j}}=2^{c_{j}K}$. Since $K$ is an integer divisible by $D$, the number $c_{j}K$ is an integer for all $j$. Hence, by introducing new variables, we may replace the constraints $x_{j}=2^{c_{j}K}$ by the corresponding LMIs of small representation size from Theorem˜5.5. The resulting SDP can be computed in polynomial time. Furthermore, it is feasible if and only if the original SDP is feasible, which proves the statement.